# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v1

@@ 1 Introduction
Mathematical models are essential for simulating real-world processes, typically mapping parameters to observable data. Estimating these parameters from real-world observations, i.e., the inverse problem, is fundamental across scientific disciplines [e.g. 10, 45, 16, 9]. However, because models never perfectly capture reality and observed data are often sparse and imprecise, parameter estimation inherently involves uncertainty. Bayesian inference offers a systematic framework for estimating parameters while incorporating uncertainty in a statistically grounded manner [e.g. 12].
@@
Markov Chain Monte Carlo (MCMC) methods are widely used for Bayesian inference to generate high-quality samples from the posterior distribution given fixed observations [13]. However, MCMC is computationally expensive and slow, rendering it impractical in scenarios where near-instant inference is required [36]. Near-instant inference is, for example, necessary in adaptive robotic control or closed-loop medical devices, where parameters such as object mass or patient sensitivity must be estimated immediately to allow accurate prediction and safe action [e.g. 23, 44, 22]. Furthermore, each new set of observations requires restarting the entire process, further increasing costs when inference is needed for multiple datasets. Multiple inference runs may be needed for ongoing adaptation as conditions evolve or more data become available. Tracking the spread of a disease such as influenza or COVID-19 is one example for this. New case data continuously arrive, and separate datasets exist for different regions, requiring repeated Bayesian updates to infer transmission rates or reproduction numbers for each location [e.g. 31, 48]. MCMC also relies on a known likelihood function that can be evaluated either analytically or numerically.
@@
Amortized Bayesian Inference (ABI), a deep-learning-based approach originating from simulation-based inference (SBI), addresses these limitations [4, 30, 21]. SBI methods are typically employed when evaluating the likelihood is infeasible, but simulations from the model are possible, allowing to learn a mapping from observed data to the posterior of the model parameters via simulated data. ABI in particular leverages generative neural networks to learn this mapping. Training data is generated through multiple model evaluations. Once trained, ABI enables near-instant posterior inference for new datasets, as the computational cost is incurred during training. Being likelihood-free makes ABI well suited for complex problems involving noisy or high-dimensional data, where evaluating the likelihood is infeasible or unreliable. However, the learned posterior only approximates the true posterior, and high accuracy requires well-designed network architectures and extensive training data.
@@
Both MCMC and ABI face limitations when the simulation model is computationally expensive, since both require a large number of model evaluations. MCMC requires many likelihood evaluations per generated posterior sample; ABI requires extensive model simulations to generate sufficient amounts of training data. As a result, if computational time is limited, obtaining a good posterior becomes infeasible for both approaches.
@@
Our goal is to enable ABI for computationally expensive models to benefit from the aforementioned advantages that ABI offers. To this end, surrogate models are the crucial tool, allowing us to reduce the computational cost of expensive simulations for training data generation. Surrogate models [42, 34], however, are only approximations (i.e., imperfect representations) of the reference simulation, and therefore not necessarily reliable. It is crucial to quantify the uncertainties associated with surrogate modeling, such as those arising from limited training data or any inherent inflexibility of the surrogate. These uncertainties then need to be propagated through the inference pipeline, ensuring consistent posterior estimation. Previous methods have integrated these uncertainties into MCMC methods with surrogate models [e.g. 20, 50, 35]. However, this requires many additional MCMC runs to incorporate the surrogate uncertainty, largely eliminating the computational advantages of surrogates.
@@
We introduce Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI), which enables ABI for computationally expensive models while accounting for the uncertainty inherent in surrogate approximations. By leveraging a surrogate, we eliminate the training error that arises from insufficient data during ABI training, as simulating from the surrogate is easy and fast. This comes at the cost of introducing a surrogate approximation error. However, unlike the ABI training error, the surrogate approximation error can be quantified and propagated, enabling reliable ABI for computationally expensive models. An illustration of the workflow, including training phases and inference, is given in Fig. 1. We demonstrate the workflow and substantiate the claimed advantages both theoretically and practically with a toy example and a real-world case study.
@@ 2.1 Amortized Bayesian Inference (ABI)
This work focuses on using ABI to infer the parameters of computationally expensive simulation models. The typical ABI procedure [30] can be divided into a training phase and an inference phase. Training includes generating training data and training a neural posterior estimator (NPE) [28, 8, 14].
@@
Training data is generated using a simulation model $M=M(\mathbf{x},\boldsymbol{\omega})$, typically a realistic first-principles model. $M$ depends on inputs $\mathbf{x}$ and parameters $\boldsymbol{\omega}$ and yields an output $\mathbf{y}$. Additionally, it can contain stochastic parts omitted in the following to simplify the notation. To learn the mapping from observations to posterior draws, a training set $D_{B}=\{(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{B}}$ of $N_{B}$ samples must first be generated. This can be performed by simulating $N_{B}$ pairs of parameters, inputs and outputs from the simulation model:
$\quad\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\quad\text{with}\quad(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\sim p(\mathbf{x},\boldsymbol{\omega}).$ (1)
@@
After creating the training dataset, the NPE, including a summary network $S_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y})$ and an inference network $I_{\boldsymbol{\varphi}}(\mathbf{s})$, must be trained. The goal of $S_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y})$ is to embed (multiple) model inputs and outputs $(\mathbf{x},\mathbf{y})$ of potentially variable length (e.g., a measurement series) in a fixed-length vector $\mathbf{s}$ that serves as input to $I_{\boldsymbol{\varphi}}(\mathbf{s})$. Then, $I_{\boldsymbol{\varphi}}(\mathbf{s})$ generates samples from the approximate posterior $q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}\mid\mathbf{s})$. Here, $\boldsymbol{\theta}$ and $\boldsymbol{\varphi}$ are learnable coefficients in $S_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y})$ and $I_{\boldsymbol{\varphi}}(\mathbf{s})$. In our case studies, we use coupling flows [29] as inference networks. Coupling flows have been empirically and theoretically shown to be expressive and allow for fast evaluation and sampling [7]. However, ABI is generally flexible in its choice for the inference network and works with any generative neural network [e.g. 1, 27, 6, 47, 38].
The summary and inference networks are trained jointly by minimizing the Kullback-Leibler (KL) divergence between the true posterior $p(\boldsymbol{\omega}\mid\mathbf{x},\mathbf{y})$ and the approximate posterior $q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}\mid\mathbf{s})$:
$L(\boldsymbol{\varphi},\boldsymbol{\theta})$ $=\mathbb{E}_{p(\mathbf{y})}[\text{KL}(p(\boldsymbol{\omega}\mid\mathbf{x},\mathbf{y})\mid q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}\mid\mathbf{s}))]\approx\sum_{i=1}^{N_{B}}(-\text{log}~{}q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}^{(i)}\mid S_{\boldsymbol{\theta}}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})))$ (2)
This expectation is approximated via the training data $D_{B}$ obtained from the (potentially expensive) simulation model [30].
@@ 2.2 Uncertainty-Aware Surrogate Modeling
A surrogate model $\widetilde{M}$ aims to approximate the behavior of a simulation model $M$ with negligible computational cost. The surrogate model is typically parametrized by learnable surrogate coefficients, denoted as $\mathbf{c}$. The simulation model output can then be approximated by
$\mathbf{y}=M(\mathbf{x},\boldsymbol{\omega})\approx\widetilde{M}_{\mathbf{c}}(\mathbf{x},\boldsymbol{\omega})=\widetilde{\mathbf{y}}.$ (3)
@@
$p(\widetilde{\mathbf{y}}_{\epsilon}\mid\widetilde{\mathbf{y}},\sigma)=p(\widetilde{\mathbf{y}}_{\epsilon}\mid\mathbf{x},\boldsymbol{\omega},\mathbf{c},\sigma),$ (5)
@@
Using sparse training data ($N_{T}\ll N_{B}$) obtained from the simulation model $D_{T}=\{(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$, with $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})$ and $(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})$ sampled from a given prior, we infer the surrogate model coefficients $\mathbf{c}$ and $\sigma$ via Bayesian inference. This enables capturing both the surrogate model’s epistemic uncertainty and the irreducible approximation error. For this purpose, we consider the likelihood (5) and the prior $p(\mathbf{c},\sigma)$. The joint posterior over the surrogate parameters and the approximation noise, given the training data, is then
$p(\mathbf{c},\sigma\mid D_{T})\propto\prod_{i=1}^{N_{T}}p(\mathbf{y}^{(i)}\mid\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma),$ (6)
@@
2.3 Surrogate-based ABI (SABI)
The quality of the NPE depends on a sufficient training budget $D_{B}$ to ensure that $p(\boldsymbol{\omega}\mid\mathbf{x},\mathbf{y})$ is well approximated via $q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}\mid S_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y}))$. However, for the expensive models considered here, generating sufficient ABI training data becomes infeasible. To mitigate this, we propose using a surrogate model of the expensive simulator to efficiently generate sufficient training data. Surrogate-based ABI (SABI) thus differs from standard ABI only in the generation of training data. Instead of using the expensive model, a (point) surrogate model with a point estimate $\overline{\mathbf{c}}$, e.g., the median of the posterior in Eq. 6 is used to generate the training data $D_{B}=\{(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\widetilde{\mathbf{y}}^{(i)})\}_{i=1}^{N_{B}}$:
$\widetilde{\mathbf{y}}^{(i)}=\widetilde{M}_{\overline{\mathbf{c}}}(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\quad\text{with }\quad(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\sim p(\mathbf{x},\boldsymbol{\omega}).$ (7)
However, in general, $\widetilde{M}_{\mathbf{c}}$ is only an approximation of the simulation model $M$. Therefore, the SABI posterior trained with surrogate data $p(\boldsymbol{\omega}\mid\mathbf{x},\widetilde{\mathbf{y}})$ will be systematically different from the true posterior $p(\boldsymbol{\omega}\mid\mathbf{x},\mathbf{y})$, as no surrogate uncertainties or errors are considered in the ABI training process. Specifically, neither the epistemic uncertainty of $\widetilde{M}_{\mathbf{c}}$, nor its irreducible approximation error is propagated. As we show, this can lead to highly inaccurate posterior approximations.

### v2

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Mathematical models are essential for simulating real-world processes, typically mapping parameters to observable data. Estimating these parameters from real-world observations, i.e., the inverse problem, is fundamental across scientific disciplines (e.g. Etz & Vandekerckhove, 2018; Von Toussaint, 2011; Hülsenbeck et al., 2001; Ellison, 2004). However, because models never perfectly capture reality and observed data are often sparse and imprecise, parameter estimation inherently involves uncertainty. Bayesian inference offers a systematic framework for estimating parameters while incorporating uncertainty in a statistically grounded manner (e.g. Gelman et al., 1995).
@@
Markov Chain Monte Carlo (MCMC) methods are widely used for Bayesian inference to generate high-quality samples from the posterior distribution given fixed observations (Gilks et al., 1995). However, MCMC is computationally expensive and slow, rendering it impractical in scenarios where near-instant inference is required (Robert et al., 2018). Near-instant inference is, for example, necessary in adaptive robotic control or closed-loop medical devices, where parameters such as object mass or patient sensitivity must be estimated immediately to allow accurate prediction and safe action (e.g. Marlier, 2024; Tasoujian et al., 2020; Malagutti et al., 2023). Furthermore, each new set of observations requires restarting the entire process, further increasing costs when inference is needed for multiple datasets. Multiple inference runs may be needed for ongoing adaptation as conditions evolve or more data become available. Tracking the spread of a disease such as influenza or COVID-19 is one example of this. New case data continuously arrive, and separate datasets exist for different regions, requiring repeated Bayesian updates to infer transmission rates or reproduction numbers for each location (e.g. Radev et al., 2021; Yang et al., 2015). MCMC also relies on a known likelihood function that can be evaluated either analytically or numerically.
@@
Amortized Bayesian Inference (ABI), a deep-learning-based approach originating from simulation-based inference (SBI), addresses these limitations (Cranmer et al., 2020; Radev et al., 2020; Lückmann et al., 2021). SBI methods are typically employed when evaluating the likelihood is infeasible, but simulations from the model are possible, allowing learning a mapping from observed data to the posterior of the model parameters via simulated data. ABI in particular leverages generative neural networks to learn this mapping. Training data is generated through multiple model evaluations. Once trained, ABI enables near-instant posterior inference for new datasets, as the computational cost is incurred during training. Being likelihood-free makes ABI well-suited for complex problems involving noisy or high-dimensional data, where evaluating the likelihood is infeasible or unreliable. However, the learned posterior only approximates the true posterior, and high accuracy requires well-designed network architectures and extensive training data.
@@
Both MCMC and ABI face limitations when the simulation model is computationally expensive, since both require a large number of model evaluations. MCMC requires many likelihood evaluations per generated posterior sample; ABI requires extensive model simulations to generate sufficient amounts of training data. As a result, if computational time is limited, obtaining a good posterior becomes infeasible for both approaches. Note that such computational effort can arise for different reasons – for instance, from a high-dimensional parameter space or from the complexity of the governing equations. In Earth sciences, for example, low-dimensional models can be computationally very expensive, particularly if they involve (coupled systems of) non-linear partial differential equations or long simulated time periods (e.g. Mohammadi et al., 2018; Hommel et al., 2015).
@@
Our goal is to enable ABI for computationally expensive models to benefit from the aforementioned advantages that ABI offers. To this end, surrogate models are the crucial tool, allowing us to reduce the computational cost of expensive simulations for training data generation. Surrogate models (Sudret, 2008; Rasmussen & Williams, 2006), however, are only approximations (i.e., imperfect representations) of the reference simulation, and therefore not necessarily reliable. It is crucial to quantify the uncertainties associated with surrogate modeling, such as those arising from limited training data or any inherent inflexibility of the surrogate. These uncertainties then need to be propagated through the inference pipeline, ensuring consistent posterior estimation. Previous methods have integrated these uncertainties into MCMC methods with surrogate models (e.g. Lückmann et al., 2018; Zhang et al., 2020; Reiser et al., 2025). However, this requires many additional MCMC runs to incorporate the surrogate uncertainty, largely eliminating the computational advantages of surrogates.
@@
We introduce Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI), which enables ABI for computationally expensive models while accounting for the uncertainty inherent in surrogate approximations. By leveraging a surrogate, we eliminate the training error that arises from insufficient data during ABI training, as simulating from the surrogate is easy and fast. This comes at the cost of introducing a surrogate approximation error. However, unlike the ABI training error, the surrogate approximation error can be quantified and propagated, enabling reliable ABI for computationally expensive models.
For this purpose, we adapt and further develop the uncertainty estimation and propagation method for surrogate modeling proposed in Reiser et al. (2025). Our main novelty is to replace the MCMC with ABI for inference. This novelty and its benefits can be viewed from two perspectives: 1) Replacing MCMC with ABI substantially improves sampling efficiency, as parallel chains for draws from the surrogate posterior are no longer required. 2) UA-SABI enables ABI training with surrogate data while accounting for the uncertainty inherent in the surrogate approximation. This significantly reduces the required runs from the expensive simulation model while retaining the amortization property and, therefore, enabling ABI for computationally expensive models.
An illustration of the workflow, including training phases and inference, is given in Fig.˜1. We demonstrate the workflow and substantiate the claimed advantages both theoretically and practically with a toy example and two real-world case studies.
@@ 2.1 Amortized Bayesian Inference (ABI)
This work focuses on using ABI to infer the parameters of computationally expensive simulation models. The typical ABI procedure (Radev et al., 2020) can be divided into a training phase and an inference phase. Training includes simulating data and then training a neural posterior estimator (NPE) on that data (Papamakarios & Murray, 2016; Durkan et al., 2019; Greenberg et al., 2019b).
@@
Training data is generated using a simulation model $M=M(\mathbf{x},\bm{\omega})$, typically a realistic first-principles model. $M$ depends on inputs $\mathbf{x}$ and parameters $\bm{\omega}$ and yields an output $\mathbf{y}$. Additionally, it can contain stochastic parts omitted in the following to simplify the notation. To learn the mapping from observations to posterior draws, a training set $D_{B}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{B}}$ of $N_{B}$ samples must first be generated. This can be performed by simulating $N_{B}$ pairs of parameters, inputs, and outputs from the simulation model:
$\quad\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\quad\text{with}\quad(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\sim p(\mathbf{x},\bm{\omega}).$ (1)
@@
After creating the training dataset, the NPE, including a summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and an inference network $I_{\bm{\varphi}}(\mathbf{s})$, must be trained. The goal of $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is to embed (multiple) model inputs and outputs $(\mathbf{x},\mathbf{y})$ of potentially variable length (e.g., a measurement series) in a fixed-length vector $\mathbf{s}$ that serves as input to $I_{\bm{\varphi}}(\mathbf{s})$. Then, $I_{\bm{\varphi}}(\mathbf{s})$ generates samples from the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$. Here, $\bm{\theta}$ and $\bm{\varphi}$ are learnable coefficients in $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and $I_{\bm{\varphi}}(\mathbf{s})$. In our case studies, we use coupling flows (Papamakarios et al., 2021) as inference networks. Coupling flows have been empirically and theoretically shown to be expressive and allow for fast evaluation and sampling (Draxler et al., 2025). However, ABI is generally flexible in its choice for the inference network and works with any generative neural network (e.g. Ardizzone et al., 2019; Padmanabha & Zabaras, 2021; Denker et al., 2021; Wildberger et al., 2023; Schmitt et al., 2024b). Thus, the summary and inference networks are modular and can be independently replaced or adapted to suit the specific problem at hand.
The summary and inference networks are trained jointly by minimizing the Kullback-Leibler (KL) divergence between the true posterior $p(\bm{\omega}\mid\mathbf{x},\mathbf{y})$ and the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$:
$L(\bm{\varphi},\bm{\theta})$ $=\mathbb{E}_{p(\mathbf{y})}[\text{KL}(p(\bm{\omega}\mid\mathbf{x},\mathbf{y})\mid q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s}))]\approx\sum_{i=1}^{N_{B}}(-\text{log}~q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid S_{\bm{\theta}}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})))$ (2)
This expectation is approximated via the training data $D_{B}$ obtained from the (potentially expensive) simulation model (Radev et al., 2020).
@@ 2.2 Uncertainty-Aware Surrogate Modeling
Uncertainty-aware surrogate modeling is introduced in and adapted from Reiser et al. (2025). A surrogate model $\widetilde{M}$ aims to approximate the behavior of a simulation model $M$ with negligible computational cost. The surrogate model is typically parametrized by learnable surrogate coefficients, denoted as $\mathbf{c}$. The simulation model output can then be approximated by
$\mathbf{y}=M(\mathbf{x},\bm{\omega})\approx\widetilde{M}_{\mathbf{c}}(\mathbf{x},\bm{\omega})=\widetilde{\mathbf{y}}$. (3)
@@
$p(\widetilde{\mathbf{y}}_{\epsilon}\mid\widetilde{\mathbf{y}},\sigma)=p(\widetilde{\mathbf{y}}_{\epsilon}\mid\mathbf{x},\bm{\omega},\mathbf{c},\sigma),$ (5)
@@
Using sparse training data ($N_{T}\ll N_{B}$) obtained from the simulation model $D_{T}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$, with $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ and $(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ sampled from a given prior, we infer the surrogate model coefficients $\mathbf{c}$ and $\sigma$ via Bayesian inference. This enables capturing both the surrogate model’s epistemic uncertainty and the irreducible approximation error. For this purpose, we consider the likelihood equation 5 and the prior $p(\mathbf{c},\sigma)$. The joint posterior over the surrogate parameters and the approximation noise, given the training data, is then
$p(\mathbf{c},\sigma\mid D_{T})\propto\prod_{i=1}^{N_{T}}p(\mathbf{y}^{(i)}\mid\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma),$ (6)
assuming an i.i.d. sampled approximation error. This assumption serves as a simple stand-in, since surrogate construction is not the focus of this work. It suffices to demonstrate that even a basic error model outperforms making no error assumption at all. In general, UA-SABI is not limited to the i.i.d. case; case-specific surrogate error models are fully compatible with the method.
The conditions under which $\widetilde{M}$ is considered optimal are implied by the assumed distributions of $\mathbf{x}$ and $\bm{\omega}$ in $D_{T}$ and the specified error model $p(\epsilon\mid\sigma)$. Sampling-based algorithms such as MCMC can be used to estimate the surrogate posterior equation 6. For the considered surrogates, this is tractable since their likelihoods are fast to evaluate, and the surrogate posterior – compared to model parameter posteriors across datasets – is estimated only once. While this formulation fully captures surrogate uncertainty, we now temporarily simplify the setup by omitting this uncertainty to introduce an uncertainty-unaware baseline method – Surrogate-based ABI. This allows us to isolate and demonstrate the benefit of propagating surrogate uncertainty through to inference.
2.3 Surrogate-based ABI (SABI)
The quality of the NPE depends on a sufficient training budget $D_{B}$ to ensure that $p(\bm{\omega}\mid\mathbf{x},\mathbf{y})$ is well approximated via $q_{\bm{\varphi}}(\bm{\omega}\mid S_{\bm{\theta}}(\mathbf{x},\mathbf{y}))$. However, for the expensive models considered here, generating sufficient ABI training data becomes infeasible. To mitigate this, we propose using a surrogate model of the expensive simulator to efficiently generate sufficient training data. Surrogate-based ABI (SABI) thus differs from standard ABI only in the generation of training data. Instead of using the expensive model, a (point) surrogate model with a point estimate $\overline{\mathbf{c}}$ is used to generate the training data $D_{B}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\widetilde{\mathbf{y}}^{(i)})\}_{i=1}^{N_{B}}$:
$\widetilde{\mathbf{y}}^{(i)}=\widetilde{M}_{\overline{\mathbf{c}}}(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\quad\text{with }\quad(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\sim p(\mathbf{x},\bm{\omega}).$ (7)
For the point surrogate, the median of the posterior in Eq.˜6 can be used as it is robust to outliers and invariant under monotonic transformations, but also the mean is a valid option.